# Algorithmic design of “smart canopy” maize architectures that maximize light use efficiency

### v1

@@ Results
The approach seamlessly extends to diverse environmental and agronomic conditions, enabling the optimization of ideotypes for various planting densities and geographic locations. We conducted independent optimization runs at three distinct maize-growing latitudes in the U.S. — Ames, IA (42°N), Thomas County, KS (39°N), and Bismarck, ND (47°N) — to assess the impact of solar position on canopy design. While trait magnitudes exhibited modest adjustments (e.g., slightly steeper upper leaves at higher latitudes), the overall architectural patterns, such as vertical stratification of leaf angles and widths, remained consistent (see Section S4 in SI for additional details). Importantly, this consistency across latitudes indicates that breeders may not need to generate location-specific canopy architectures for commercial production, thereby simplifying the deployment of improved ideotypes at scale.
@@ Discussion
Beyond azimuth, the optimization favored erect upper leaves and tapered widths—traits known to reduce mutual shading and promote light transmission to lower canopy strata. The preference for upright upper leaves mirrors phenotypes in elite commercial hybrids (e.g., Pioneer P1151AM™) and aligns with genetic studies on liguleless1 and LAC1, which demonstrate that reduced leaf inclination at the canopy top can enhance light penetration and photosynthetic efficiency tian2024maize. These convergences validate the biological realism of the optimization and reinforce empirical evidence that targeted control of leaf angle can improve canopy light-use efficiency. To validate this, future work should involve phenotyping of genetic lines engineered or selected to approximate the optimized trait combinations, followed by field trials under representative agronomic conditions.
@@
From a breeding perspective, the ability to algorithmically explore and rank thousands of virtual trait combinations provides a powerful complement to traditional field selection. Breeding pipelines are often constrained by long cycle times, environmental variability, and the difficulty of disentangling genetic linkage and pleiotropy. In contrast, computational ideotype design allows breeders to pre-screen architectural configurations under controlled virtual environments, identify promising multi-trait combinations, and focus empirical testing on only the most promising candidates. This targeted approach can reduce trial-and-error in the field, guide marker-assisted selection toward architecture-associated loci, and prioritize gene-editing interventions for traits like leaf angle and azimuth that have clear mechanistic links to light capture.
@@
A major strength of our pipeline is the integration of an evolutionary algorithm with a physics-based reverse ray-tracing model (Helios bailey2019helios) that accounts for dynamic solar position, mutual shading, and multiple scattering with mesh-level geometric precision. This enables evaluation of trait combinations under realistic full-diurnal light conditions, avoiding the oversimplifications common in empirical light-distribution models. By evaluating performance at the R1 stage abendroth2011corn, we target the critical phenological window where canopy architecture is fully established and grain filling begins. This isolates the effects of geometry from growth dynamics, providing a ’static’ blueprint that can represent the cumulative architectural goal for breeding programs. Further, the simulation workflow used here follows a previously documented, field-referenced pipeline for estimating canopy PAR saleem2025accessing, and the trait bounds and organ definitions used to build our parametric plants are drawn from segmented, field-derived 3D maize reconstructions with explicit leaf-level parameterization MaizeField3D. Together, these links to empirical data support both the physiological plausibility and reproducibility of the in silico ideotypes.
@@ Conclusion
This study introduces an end-to-end framework for computational smart canopy ideotype design, combining evolutionary algorithm, mechanistic 3D plant modeling, and physics-based light simulation to identify maize architectures that maximize canopy-level light interception. The approach recovered known advantageous traits (e.g., erect upper leaves) and revealed underappreciated configurations (e.g., tiled azimuthal arrangements) that challenge conventional assumptions about optimal plant form. Results were robust across latitudes and planting densities (SI), and the pipeline is grounded in empirically derived organ-level geometry and a validated radiative-transfer configuration, strengthening confidence in the simulated gains.

### v2

@@ 1 Introduction
The approach seamlessly extends to diverse environmental and agronomic conditions, enabling the optimization of ideotypes for various planting densities and geographic locations. We conducted independent optimization runs at five distinct latitudes spanning the southmost to northmost limits of commercial maize production above the equator, including three maize-growing sites in the U.S. — Ames, IA (42N), Thomas County, KS (39N), and Bismarck, ND (47N) — as well as Barinas, Venezuela (7N) representing the southern extreme and Peace River, Alberta, Canada (56N) representing the northern extreme to assess the impact of solar position on canopy design. While trait magnitudes exhibited modest adjustments (e.g., slightly steeper upper leaves at higher latitudes), the overall architectural patterns, such as vertical stratification of leaf angles and widths, remained consistent (see Section S4 in SI for additional details). Importantly, this consistency across latitudes indicates that breeders may not need to generate location-specific canopy architectures for commercial production, thereby simplifying the deployment of improved ideotypes at scale.
@@ Discussion
Beyond azimuth, the optimization favored erect upper leaves and tapered widths—traits known to reduce mutual shading and promote light transmission to lower canopy strata. The preference for upright upper leaves mirrors phenotypes in elite commercial hybrids (e.g., Pioneer P1151AM™) and aligns with genetic studies on liguleless1 and LAC1, which demonstrate that reduced leaf vertical angle at the canopy top can enhance light penetration and photosynthetic efficiency [tian2024maize]. These convergences validate the biological realism of the optimization and reinforce empirical evidence that targeted control of leaf angle can improve canopy light-use efficiency. To validate this, future work should involve phenotyping of genetic lines engineered or selected to approximate the optimized trait combinations, followed by field trials under representative agronomic conditions.
@@
From a breeding perspective, the ability to algorithmically explore and rank thousands of virtual trait combinations provides a powerful complement to traditional field selection. Breeding pipelines are often constrained by long cycle times, environmental variability, and the difficulty of disentangling genetic linkage and pleiotropy. In contrast, computational ideotype design allows breeders to pre-screen architectural configurations under controlled virtual environments, identify promising multi-trait combinations, and focus empirical testing on only the most promising candidates. This targeted approach can reduce trial-and-error in the field, guide marker-assisted selection toward architecture-associated loci, and prioritize gene-editing interventions for traits like leaf angle and azimuthal angle that have clear mechanistic links to light capture.
@@
A major strength of our pipeline is the integration of an evolutionary algorithm with a physics-based reverse ray-tracing model (Helios [bailey2019helios]) that accounts for dynamic solar position, mutual shading, and multiple scattering with mesh-level geometric precision. This enables evaluation of trait combinations under realistic full-diurnal light conditions, avoiding the oversimplifications common in empirical light-distribution models. By evaluating performance at the R1 stage [abendroth2011corn], we target the critical phenological window where canopy architecture is fully established and grain filling begins. This isolates the effects of geometry from growth dynamics, providing a ’static’ blueprint that can represent the cumulative architectural goal for breeding programs. Further, the simulation workflow used here follows a previously documented, field-referenced pipeline for estimating canopy PAR [saleem2025accessing], and the trait bounds and organ definitions used to build our parametric plants are drawn from segmented, field-derived 3D maize reconstructions with explicit leaf-level parameterization [MaizeField3D]. Together, these links to empirical data support both the physiological plausibility and reproducibility of the in silico ideotypes.
@@ Conclusion
This study introduces an end-to-end framework for computational smart canopy ideotype design, combining evolutionary algorithm, mechanistic 3D plant modeling, and physics-based light simulation to identify maize architectures that maximize canopy-level light interception. The approach recovered known advantageous traits (e.g., erect upper leaves) and revealed underappreciated configurations (e.g., quasi-radial azimuthal arrangements) that challenge conventional assumptions about optimal plant form. Results were robust across latitudes and planting densities (Section S4 and S5 in SI), and the pipeline is grounded in empirically derived organ-level geometry and a validated radiative-transfer configuration, strengthening confidence in the simulated gains.